# MM-Vet v2: A Challenging Benchmark to Evaluate Large Multimodal Models for Integrated Capabilities

### v1

@@ 3 Experiments
OpenFlamingo-9B [4, 6] 24.8 19.1 11.0 10.5 13.0 18.3 10.2 3.1 17.6$\pm$0.2
Otter-9B [13] 24.7 25.1 16.5 13.0 19.5 23.4 9.9 6.8 23.2$\pm$0.1
LLaVA-v1.5-7B [17] 34.2 30.4 21.5 21.6 25.4 23.5 9.5 6.4 28.3$\pm$0.2
LLaVA-v1.5-13B [17] 39.2 34.8 29.2 28.6 29.1 29.3 17.8 6.2 33.2$\pm$0.1
CogAgent-Chat [12] 40.5 33.6 29.2 34.9 27.3 33.6 24.7 6.2 34.7$\pm$0.2
Emu2-Chat [25] 45.5 37.9 28.8 35.1 35.1 37.7 18.7 8.3 38.0$\pm$0.1
IXC2-VL-7B [9] 52.3 40.7 37.0 42.1 37.0 39.0 5.2 37.7 42.5$\pm$0.3
CogVLM-Chat [27] 52.6 46.3 40.7 42.2 40.3 47.8 33.8 8.2 45.1$\pm$0.2
InternVL-Chat-V1-2 [8] 49.7 46.1 40.8 43.1 41.4 41.8 25.1 9.4 45.5$\pm$0.1
LLaVA-NeXT-34B [17] 60.2 49.3 48.9 53.2 48.3 49.6 18.5 37.3 50.9$\pm$0.1
InternVL-Chat-V1-5 [8] 62.8 52.0 48.9 51.7 49.3 47.9 37.6 17.6 51.5$\pm$0.2
InternVL2-40B [3] 61.8 63.6 63.9 64.3 60.0 60.5 48.5 45.0 63.8$\pm$0.2
InternVL2-Llama3-76B [3] 64.4 67.0 68.5 71.2 65.2 62.0 59.4 58.8 68.4$\pm$0.3
@@
Qwen-VL-Max [7] 66.6 51.7 51.1 60.2 49.0 52.2 27.3 58.3 55.8$\pm$0.2
Gemini Pro Vision [26] 63.1 54.3 50.8 61.9 55.8 50.7 45.4 46.3 57.2$\pm$0.2
GPT-4V [20] 67.7 63.1 67.1 74.7 65.8 53.1 62.2 72.6 66.3$\pm$0.2
Gemini 1.5 Pro [22] 65.8 64.4 64.7 75.6 65.9 56.2 63.9 65.3 66.9$\pm$0.2
GPT-4o [2] 69.3 67.5 70.5 79.1 68.8 63.4 74.3 79.4 71.0$\pm$0.2
Claude 3.5 Sonnet [1] 74.2 69.2 70.8 79.6 68.8 65.8 67.7 70.3 71.8$\pm$0.2
@@
Model Gen Know Rec Rec Gen Ocr Rec Spat Gen Rec Ocr Rec Spat OCR Spat Gen Rec Spat Math Ocr Spat Know Rec Know Rec Spat Gen Rec Seq Gen Ocr Gen Know Ocr Rec Ocr Rec Spat Math OCR Total
OpenFlamingo-9B [4, 6] 24.8 14.9 42.7 6.1 18.3 20.2 20.0 14.6 8.1 6.7 40.0 13.3 34.6 0.0 2.8 16.7 17.6$\pm$0.2
Otter-9B [13] 24.7 20.5 48.2 5.0 38.8 27.2 49.6 17.3 10.9 14.5 20.7 8.4 55.0 0.0 17.7 12.5 23.2$\pm$0.1
LLaVA-v1.5-7B [17] 23.7 64.2 24.5 36.5 38.4 60.4 19.6 10.9 13.3 13.3 33.6 2.9 0.5 24.6 12.5 0.0 28.3$\pm$0.2
LLaVA-v1.5-13B [17] 26.5 60.7 33.9 55.5 49.9 44.8 28.5 12.8 13.3 30.0 39.0 16.9 0.0 43.4 20.8 0.0 33.2$\pm$0.1
CogAgent-Chat [12] 30.1 59.7 29.3 33.6 64.9 40.4 44.2 3.5 13.3 33.3 33.1 26.6 26.5 60.2 8.3 0.0 34.7$\pm$0.2
Emu2-Chat [25] 34.6 67.1 38.8 28.6 64.4 63.7 48.0 5.2 13.3 49.3 49.6 21.7 5.7 38.3 8.3 0.0 38.0$\pm$0.1
IXC2-VL-7B [9] 40.9 66.9 40.4 60.9 71.1 67.0 48.1 20.4 46.0 14.7 36.3 3.6 13.8 56.3 0.0 36.4 42.5$\pm$0.3
CogVLM-Chat [27] 44.4 72.2 36.9 46.1 53.6 58.1 55.7 18.7 10.8 32.7 53.6 32.7 22.2 61.2 33.3 9.1 45.1$\pm$0.2
LLaVA-NeXT-34B [17] 53.2 66.7 57.0 55.1 71.6 67.7 65.0 44.4 46.7 26.3 34.6 10.0 32.3 73.5 16.7 34.5 50.9$\pm$0.1
InternVL-Chat-V1-5 [8] 47.5 72.1 59.2 56.6 69.1 67.7 69.2 37.7 23.6 37.3 31.0 42.7 22.9 70.8 25.0 0.0 51.5$\pm$0.2
InternVL2-40B 61.8 57.8 77.7 70.0 75.0 81.6 77.8 62.3 59.9 53.3 40.4 57.9 61.4 54.6 86.5 25.0 63.8$\pm$0.2
InternVL2-Llama3-76B [3] 64.4 62.4 80.4 74.5 77.7 75.2 81.0 79.5 60.5 66.7 43.3 67.7 40.0 74.8 89.1 25.0 68.4$\pm$0.3
@@
Qwen-VL-Max [7] 52.6 71.2 53.5 59.2 76.8 67.0 76.7 41.9 66.0 60.0 23.6 33.4 68.6 79.1 8.3 63.6 55.8$\pm$0.2
Gemini Pro Vision [26] 45.4 67.3 51.3 59.1 83.8 75.3 76.7 48.6 52.0 64.7 42.9 39.9 50.6 67.4 33.3 51.8 57.2$\pm$0.2
GPT-4V [20] 56.4 71.0 82.4 77.6 85.5 65.5 80.4 62.6 86.0 32.7 25.0 62.7 64.2 80.0 41.7 63.6 66.3$\pm$0.2
Gemini 1.5 Pro [22] 54.3 71.0 79.0 70.3 86.6 80.0 73.1 44.4 73.3 64.7 42.1 77.1 78.9 76.0 41.7 63.6 66.9$\pm$0.2
GPT-4o [2] 62.4 71.1 79.1 79.5 89.9 67.9 82.5 54.1 99.3 47.2 51.9 75.7 73.5 86.2 25.0 63.6 71.0$\pm$0.2
Claude 3.5 Sonnet [1] 65.2 79.7 79.2 68.5 83.8 72.6 86.5 48.4 86.7 50.0 43.6 81.7 80.6 83.8 66.7 54.5 71.8$\pm$0.2
@@ 3.1 Experiment settings
We evaluate two types of LMMs on our MM-Vet v2: (1) open-weight LMMs including OpenFlamingo [4, 6], Otter [13], LLaVA [17, 16], CogAgent [12], Emu2-Chat [25], InternLM-XComposer2 (IXC2) [9], CogVLM [27] and InternVL [8]; (2) Closed source models including Claude [5, 1], Qwen-VL Max [7], Gemini [26, 22], GPT-4V [20] and GPT-4o [2]).

### v2

@@ 3 Experiments
OpenFlamingo-9B [4, 6] 24.8 19.1 11.0 11.2 13.0 18.4 10.2 2.9 17.6$\pm$0.2
Otter-9B [13] 24.7 25.1 16.5 13.7 19.5 23.5 9.9 6.4 23.2$\pm$0.1
LLaVA-v1.5-7B [17] 34.2 30.4 21.5 21.7 25.4 23.6 9.5 6.0 28.3$\pm$0.2
LLaVA-v1.5-13B [17] 39.2 34.8 29.2 28.7 29.1 29.4 17.8 8.8 33.2$\pm$0.1
CogAgent-Chat [12] 40.5 33.6 29.2 35.3 27.3 33.9 24.7 5.9 34.7$\pm$0.2
Emu2-Chat [25] 45.5 37.9 28.8 35.1 35.1 37.9 18.7 10.8 38.0$\pm$0.1
IXC2-VL-7B [9] 52.3 40.7 37.0 42.5 37.0 39.3 5.2 36.1 42.5$\pm$0.3
CogVLM-Chat [27] 52.6 46.3 40.7 41.6 40.3 47.5 33.8 7.7 45.1$\pm$0.2
InternVL-Chat-V1-2 [8] 49.7 46.1 40.8 43.8 41.4 42.1 25.1 11.8 45.5$\pm$0.1
LLaVA-NeXT-34B [17] 60.2 49.3 48.9 52.8 48.3 49.9 18.5 35.1 50.9$\pm$0.1
InternVL-Chat-V1-5 [8] 62.8 52.0 48.9 51.0 49.3 48.2 37.6 16.5 51.5$\pm$0.2
InternVL2-40B [3] 61.8 63.6 63.9 64.3 60.0 60.9 48.5 45.3 63.8$\pm$0.2
InternVL2-Llama3-76B [3] 64.4 67.0 68.5 70.6 65.2 62.4 59.4 55.4 68.4$\pm$0.3
@@
Qwen-VL-Max [7] 66.6 51.7 51.1 60.2 49.0 52.5 27.3 60.2 55.8$\pm$0.2
Gemini Pro Vision [26] 63.1 54.3 50.8 61.5 55.8 50.4 45.4 43.5 57.2$\pm$0.2
GPT-4V [20] 67.7 63.1 67.1 73.6 65.8 53.4 62.2 68.3 66.3$\pm$0.2
Gemini 1.5 Pro [22] 65.8 64.4 64.7 75.1 65.9 56.6 63.9 61.5 66.9$\pm$0.2
GPT-4o [2] 69.3 67.5 70.5 78.0 68.8 63.8 74.3 77.6 71.0$\pm$0.2
Claude 3.5 Sonnet [1] 74.2 69.2 70.8 78.9 68.8 65.6 67.7 69.1 71.8$\pm$0.2
@@
Model Rec Gen Know Rec Rec Gen OCR Spat Rec Gen OCR Rec Spat OCR Spat Rec Gen Spat OCR Spat Math Rec Gen Seq Gen OCR Rec Know Rec Gen OCR Know Rec OCR Spat OCR Math Rec Spat Know Total
OpenFlamingo-9B [4, 6] 14.9 42.7 6.1 18.3 20.2 19.3 14.6 8.1 6.7 13.3 0.0 46.2 2.8 16.7 0.0 30.4 17.6$\pm$0.2
Otter-9B [13] 20.5 48.2 5.0 38.8 27.2 47.8 17.3 10.9 14.5 8.4 0.0 23.8 17.7 12.5 0.0 58.0 23.2$\pm$0.1
LLaVA-v1.5-7B [17] 23.7 64.2 24.5 36.5 38.4 58.2 19.6 10.9 13.3 2.9 0.5 15.4 24.6 12.5 0.0 37.0 28.3$\pm$0.2
LLaVA-v1.5-13B [17] 26.5 60.7 33.9 55.5 49.9 43.2 28.5 12.8 13.3 16.9 0.0 26.9 43.4 20.8 0.0 44.6 33.2$\pm$0.1
CogAgent-Chat [12] 30.1 59.7 29.3 33.6 64.9 38.9 44.2 3.5 13.3 26.6 26.5 38.5 60.2 8.3 0.0 26.4 34.7$\pm$0.2
Emu2-Chat [25] 34.6 67.1 38.8 28.6 64.4 61.4 48.0 5.2 13.3 21.7 5.7 49.2 38.3 8.3 0.0 59.4 38.0$\pm$0.1
IXC2-VL-7B [9] 40.9 66.9 40.4 60.9 71.1 64.6 48.1 20.4 46.0 3.6 13.8 15.4 56.3 0.0 36.4 30.8 42.5$\pm$0.3
CogVLM-Chat [27] 44.4 72.2 36.9 46.1 53.6 59.6 55.7 18.7 10.8 32.7 22.2 37.7 61.2 33.3 9.1 65.0 45.1$\pm$0.2
InternVL-Chat-V1-2 [8] 39.6 69.9 46.9 57.9 75.7 71.8 48.1 22.2 13.3 32.3 18.6 29.2 64.2 16.7 9.1 35.2 45.5$\pm$0.1
LLaVA-NeXT-34B [17] 53.2 66.7 57.0 55.1 71.6 65.3 65.0 44.4 46.7 10.0 32.3 30.3 73.5 16.7 34.5 41.6 50.9$\pm$0.1
InternVL-Chat-V1-5 [8] 47.5 72.1 59.2 56.6 69.1 65.3 69.2 37.7 23.6 42.7 22.9 43.1 70.8 25.0 0.0 43.4 51.5$\pm$0.2
InternVL2-40B 57.8 77.7 70.0 75.0 81.6 75.0 62.3 59.9 53.3 57.9 54.6 38.9 86.5 25.0 40.9 66.8 63.8$\pm$0.2
InternVL2-Llama3-76B [3] 62.4 80.4 74.5 77.7 75.2 78.1 79.5 60.5 66.7 67.7 74.8 50.0 89.1 25.0 63.6 47.4 68.4$\pm$0.3
@@
Qwen-VL-Max [7] 52.6 71.2 53.5 59.2 76.8 64.6 76.7 41.9 66.0 33.4 68.6 55.4 79.1 8.3 63.6 15.0 55.8$\pm$0.2
Gemini Pro Vision [26] 45.4 67.3 51.3 59.1 83.8 76.1 76.7 48.6 52.0 39.9 50.6 74.6 67.4 33.3 51.8 40.0 57.2$\pm$0.2
GPT-4V [20] 56.4 71.0 82.4 77.6 85.5 63.1 80.4 62.6 86.0 62.7 64.2 37.7 80.0 41.7 63.6 35.0 66.3$\pm$0.2
Gemini 1.5 Pro [22] 54.3 71.0 79.0 70.3 86.6 77.1 73.1 44.4 73.3 77.1 78.9 74.6 76.0 41.7 63.6 45.0 66.9$\pm$0.2
GPT-4o [2] 62.4 71.1 79.1 79.5 89.9 65.5 82.5 54.1 99.3 75.7 73.5 46.8 86.2 25.0 63.6 72.6 71.0$\pm$0.2
Claude 3.5 Sonnet [1] 65.2 79.7 79.2 68.5 83.8 73.6 86.5 48.4 86.7 81.7 80.6 50.0 83.8 66.7 54.5 41.0 71.8$\pm$0.2
@@ 3.1 Experiment settings
We evaluate two types of LMMs on our MM-Vet v2: (1) open-weight LMMs including OpenFlamingo [4, 6], Otter [13], LLaVA [17, 16], CogAgent [12], Emu2-Chat [25], InternLM-XComposer2 (IXC2) [9], CogVLM [27] and InternVL [8]; (2) Closed source models including Claude [5, 1], Qwen-VL Max [7], Gemini [26, 22], GPT-4V [20] and GPT-4o [2].